\begin{thebibliography}{10}

\bibitem{Ben-Artzi/Gohberg:94}
A.~Ben-Artzi and I.~Gohberg.
\newblock {Orthogonal polynomials over Hilbert modules}.
\newblock In A.~Feintuch et~al., editor, {\em {Nonselfadjoint operators and
  related topics. Workshop on Operator theory and its applications, Beersheva,
  Israel, February 24-28, 1992}}, volume~{\bf 73} of {\em Basel,
  Birkh\"auser-Verlag, {\it Oper. Theory, Adv. Appl.}}, (1994).
\newblock 96-126.

\bibitem{Biedenharn/Horwitz:84}
L.~C. Biedenharn and L.~P. Horwitz.
\newblock {Quaternion quantum mechanics: second quantisation and gauge fields}.
\newblock {\em Ann. Phys. (USA)}, {\bf 157}, (1984).
\newblock 432-488.

\bibitem{Biedenharn/Horwitz:85}
L.~C. Biedenharn and L.~P. Horwitz.
\newblock {Erratum: Quaternion quantum mechanics: second quantisation and gauge
  fields}.
\newblock {\em Ann. Phys. (USA)}, {\bf 159}, (1985).
\newblock 481.

\bibitem{Bultheel:80}
A.~Bultheel.
\newblock {Orthogonal matrix functions related to the multivariable
  Nevanlinna-Pick problem}.
\newblock {\em Bull. Soc. Math. Belg., Ser. B}, {\bf 32}, (1980).
\newblock 149-170.

\bibitem{Bultheel:82}
A.~Bultheel.
\newblock {Inequalities in Hilbert modules of matrix-valued functions}.
\newblock {\em Proc. Amer. Math. Soc.}, {\bf 85}, (1982).
\newblock 369-372.

\bibitem{Cnops:92}
J.~Cnops.
\newblock {A Gram-Schmidt method in Hilbert modules}.
\newblock In {\em Clifford Algebras and Their Apllications in Mathematical
  Physics (Montpellier, 1989), p. 193-203, {\it Fund. Theories Physics} {\bf
  47}, Kluwer Acad. Publ., Dordrecht}, 1992.

\bibitem{Delanghe:79}
R.~Delanghe.
\newblock {On the extension of linear functionals in modules over an
  H*-algebra}.
\newblock {\em Simon Stevin}, {\bf 53}(no.3), (1979).
\newblock 201-209.

\bibitem{Delanghe/Brackx:78}
R.~Delanghe and F.~Brackx.
\newblock {Hypercomplex function theory and Hilbert modules with reproducing
  kernel}.
\newblock {\em Proc. London Math. Soc.}, {\bf 37}, (1978).
\newblock 545-576.

\bibitem{Ellis/Gohberg/Lay:95}
R.~L. Ellis, I~Gohberg, and D.~C. Lay.
\newblock {Infinite analogues of block Toeplitz matrices and related orthogonal
  functions}.
\newblock {\em Integral Equations Oper. Theory}, {\bf 22}, (1995).
\newblock 375-419.

\bibitem{Fuge:95}
H.~Fuge.
\newblock {Einige Aussagen \"uber beschr\"ankte Operatoren in Hilbertmoduln}.
\newblock Master's thesis, Univ. Leipzig, 1995.

\bibitem{Cabrera:91}
M.~Cabrera Garc\'{\i}a.
\newblock {Hilbert modules on H*-algebras and applications(span.)}.
\newblock In {\em Analysis}. 15th Port.-Span. Meeting Math., Evora/Port., 1990,
  v.~II, 1991.
\newblock 297-302.

\bibitem{Cabrera/ElMarrakchi/Martinez/Rodriguez:93}
M.~Cabrera Garc\'{\i}a, A.~El Marrakchi, J.~Mart\'{\i}nez Moreno, and
  A.~Rodr\'{\i}guez Palacios.
\newblock {Bounded differential operators on Hilbert modules and derivations of
  structurable H*-algebras}.
\newblock {\em Commun. Algebra}, {\bf 21}, (1993).
\newblock 2905-2945.

\bibitem{Cabrera/Martinez/Rodriguez:89/2}
M.~Cabrera Garc\'{\i}a, J.~Mart\'{\i}nez Moreno, and A.~Rodr\'{\i}guez
  Palacios.
\newblock {Structurable H*-algebras}.
\newblock preprint 1989.
\newblock Univ. de Granada, Spain.

\bibitem{Cabrera/Martinez/Rodriguez:92}
M.~Cabrera Garc\'{\i}a, J.~Mart\'{\i}nez Moreno, and A.~Rodr\'{\i}guez
  Palacios.
\newblock {Hilbert modules over H*-algebras in relation with Hilbert ternary
  rings}.
\newblock In {\em Nonassociative algebraic models}, 1992.
\newblock Proc. Workshop, Zaragoza, Spain, 1989, 33-44.

\bibitem{Cabrera/Martinez/Rodriguez:95}
M.~Cabrera Garc\'{\i}a, J.~Mart\'{\i}nez Moreno, and A.~Rodr\'{\i}guez
  Palacios.
\newblock {Hilbert modules revised: Orthonormal bases and Hilbert-Schnidt
  operators}.
\newblock {\em Glasgow Math. J.}, {\bf 37}, (1995).
\newblock 45-54.

\bibitem{Giellis:72}
G.~R. Giellis.
\newblock {A characterization of Hilbert modules}.
\newblock {\em Proc. Amer. Math. Soc.}, {\bf 36}, (1972).
\newblock 440-442.

\bibitem{Goldstine/Horwitz:66}
H.~H. Goldstine and L.~P. Horwitz.
\newblock {Hilbert space with non-associative scalars}.
\newblock {\em Math. Ann.}, {\bf 164}, (1966).
\newblock 291-316.

\bibitem{Kokschal:95}
A.~Kokschal.
\newblock {Einige Grundz\"uge der Theorie der Hilbertmoduln}.
\newblock Master's thesis, Univ. Leipzig, 1995.
\newblock 379 pp.

\bibitem{Kulkarni:94}
S.~H. Kulkarni.
\newblock {Representation of a real B*-algebra on a quaternionic Hilbert
  space}.
\newblock {\em Proc. Amer. Math. Soc.}, {\bf 121}, (1994).
\newblock 505-509.

\bibitem{Masani:59}
P.~R. Masani.
\newblock {Cram\'ers theorem on monotone matrix-valued functions and the Wold
  decomposition}.
\newblock In U.~Grenande, editor, {\em {Probability and Statistics - The Harald
  Cram\'er Volume}}. Almquist \& Wiksell, Stockholm, 1959.
\newblock 175-189.

\bibitem{Wiener/Masani:60}
P.~R. Masani.
\newblock {The prediction theory of multivariable stochastic processes, III.
  Unbounded spectral densities}.
\newblock {\em Acta Math.}, {\bf 104}, (1960).
\newblock 141-162.

\bibitem{Masani:62}
P.~R. Masani.
\newblock {Shift invariant spaces and prediction theory}.
\newblock {\em Acta Math.}, {\bf 107}, (1962).
\newblock 275-290.

\bibitem{Masani:66}
P.~R. Masani.
\newblock {Recent trends in multivariate prediction theory}.
\newblock In P.~R. Krishnaiah, editor, {\em {Multivariate Analysis, Proc. Int.
  Symp., Dayton, Ohio, June 1965}}. Academic Press, New York, 1966.
\newblock 351-381.

\bibitem{Molnar:90}
L.~Moln\'ar.
\newblock {Reproducing kernel Hilbert $A$-modules}.
\newblock {\em Glasnik Matemati{\v{c}}ki}, {\bf 25(45)}, (1990).
\newblock 335-345.

\bibitem{Molnar:91}
L.~Moln\'ar.
\newblock {A note on Saworotnow's representation theorem on positive definite
  functions}.
\newblock {\em Houston J. Math.}, {\bf 17}, (1991).
\newblock 89-99.

\bibitem{Molnar:92}
L.~Moln\'ar.
\newblock {A note on the strong Schwarz inequality in Hilbert $A$-modules}.
\newblock {\em Publ. math. Debrecen}, {\bf 40}, (1992).
\newblock 323-325.

\bibitem{Molnar:92/3}
L.~Moln\'ar.
\newblock {Modular bases in a Hilbert $A$-module}.
\newblock {\em Czech. Math. J.}, {\bf 42}, (1992).
\newblock 649-656.

\bibitem{Molnar:92/2}
L.~Moln\'ar.
\newblock {Representations of the trace class of an H*-algebra}.
\newblock {\em Proc. Amer. Math. Soc.}, {\bf 115}, (1992).
\newblock 167-170.

\bibitem{Molnar:93/1}
L.~Moln\'ar.
\newblock {On A-linear operators on a Hilbert $A$-module}.
\newblock {\em Period. Math. Hungar.}, {\bf 26}, (1993).
\newblock 219-222.

\bibitem{Molnar:93/2}
L.~Moln\'ar.
\newblock {On Saworotnow's Hilbert $A$-modules}.
\newblock {\em Glas. Mat.}, {\bf 28}, (1993).
\newblock 259-287.

\bibitem{Razon/Horwitz:91/2}
A.~Razon and L.~P. Horwitz.
\newblock {Projection operators and states in the tensor product of quaternion
  Hilbert modules}.
\newblock {\em Acta Appl. Math.}, {\bf 24}, (1991).
\newblock 179-194.

\bibitem{Razon/Horwitz:91/1}
A.~Razon and L.~P. Horwitz.
\newblock {Tensor products of quaternion Hilbert modules}.
\newblock {\em Acta Appl. Math.}, {\bf 24}, (1991).
\newblock 141-178.

\bibitem{Razon/Horwitz:92}
A.~Razon and L.~P. Horwitz.
\newblock {Uniqueness of the scalar product in the tensor product of quaternion
  Hilbert modules}.
\newblock {\em J. Math. Phys.}, {\bf 33}, (1992).
\newblock 3098-3104.

\bibitem{Razon/Horwitz/Biedenharn:89}
A.~Razon, L.~P. Horwitz, and L.~C. Biedenharn.
\newblock {On a basic theorem of quaternionic modules}.
\newblock {\em J. Math. Phys.}, {\bf 30}, (1989).
\newblock 59.

\bibitem{Rosenberg:64}
M.~Rosenberg.
\newblock {The square-integrability of matrix-valued functions with respect to
  a non-negative hermitian measure}.
\newblock {\em Duke Math. J.}, {\bf 31}, (1964).
\newblock 291-298.

\bibitem{Salehi:65}
H.~Salehi.
\newblock {\em {The prediction theory of multivariate stochastic processes with
  continuous time}}.
\newblock PhD thesis, Indiana Univ., Bloomington, U.S.A., 1965.

\bibitem{Salehi:66}
H.~Salehi.
\newblock {A factorization algorithm for q$\times$q matrix-valued functions on
  the real line R}.
\newblock {\em Trans. Amer. Math. Soc.}, {\bf 124}, (1966).
\newblock 468-470.

\bibitem{Salehi:67}
H.~Salehi.
\newblock {On the growth of a q-variate stationary stochastic process}.
\newblock {\em Z. Wahrscheinlichkeitstheorie und verw. Gebiete}, {\bf 8},
  (1967).
\newblock 140-147.

\bibitem{Saworotnow:68}
P.~P. Saworotnow.
\newblock {A generalized Hilbert space}.
\newblock {\em Duke Math. J.}, {\bf 35}, (1968).
\newblock 191-197.

\bibitem{Saworotnow:70}
P.~P. Saworotnow.
\newblock {Representation of a topological group on a Hilbert module}.
\newblock {\em Duke Math. J.}, {\bf 37}, (1970).
\newblock 145-150.

\bibitem{Saworotnow:80}
P.~P. Saworotnow.
\newblock {Linear spaces with an H*-algebra-valued inner product}.
\newblock {\em Trans. Amer. Math. Soc,}, {\bf 262}, (1980).
\newblock 543-549.

\bibitem{Saworotnow:81}
P.~P. Saworotnow.
\newblock {Irreducible representations of a Banach algebra on a Hilbert
  module}.
\newblock {\em Houston J. Math.}, {\bf 7}, (1981).
\newblock 275-281.

\bibitem{Saworotnow:82}
P.~P. Saworotnow.
\newblock {Errata to: Linear spaces with an H*-algebra-valued inner product}.
\newblock {\em Trans. Amer. Math. Soc,}, {\bf 270}, (1982).
\newblock 349.

\bibitem{Saworotnow:83}
P.~P. Saworotnow.
\newblock {Generalized positive definite functions and stationary processes}.
\newblock In V.~Mandrekar and H.~Salehi, editors, {\em {Prediction Theory and
  Harmonic Analysis, The Pesi Masani volume}}. North-Holland Publishing Comp.,
  Amsterdam, 1983.
\newblock 329-344.

\bibitem{Saworotnow:85}
P.~P. Saworotnow.
\newblock {Diagonalization of a self-adjoint operator acting on a Hilbert
  module}.
\newblock {\em Int. J. Math. Math. Sci.}, {\bf 8}, (1985).
\newblock 669-675.

\bibitem{Saworotnow:91}
P.~P. Saworotnow.
\newblock {Diagonalization of unbounded self-adjoint operators acting on a
  Hilbert module}.
\newblock {\em Glas. Mat., III.~Series}, {\bf 26(46)}, (1991).
\newblock 89-95.

\bibitem{Sen:82}
D.~K. Sen.
\newblock {On operators on Hilbert spaces over H*-algebras}.
\newblock {\em An. Stiint. Univ. Al. I. Cuza Iasi, N. Ser., Sect. Ia}, {\bf
  28}, no.~2, (1982).
\newblock 29-33.

\bibitem{Sharma/Coulson:87}
C.~S. Sharma and T.~J. Coulson.
\newblock {Spectral theory for unitary operators on a quaternionic Hilbert
  space}.
\newblock {\em J. Math. Phys.}, {\bf 28}, (1987).
\newblock 1941-1945.

\bibitem{Smith:73}
J.~F. Smith.
\newblock {The p-classes of a Hilbert module}.
\newblock {\em Proc. Amer. Math. Soc.}, {\bf 36}, (1973).
\newblock 428-434.

\bibitem{Smith:74}
J.~F. Smith.
\newblock {The structure of Hilbert modules}.
\newblock {\em J. London Math. Soc.}, {\bf 8}, (1974).
\newblock 741-749.

\bibitem{Soffer/Horwitz:83}
A.~Soffer and L.~P. Horwitz.
\newblock {B*-algebra representations in a quaternionic Hilbert module}.
\newblock {\em J. Math. Phys.}, {\bf 24}, (1983).
\newblock 2780-2782.

\bibitem{Truong-Van:81}
B.~Truong-Van.
\newblock {Une g\'en\'eralisation du th\'eor\`eme de Kolmogorov-Aronszajn.
  Processus V-born\'es q-dimensionelles: domaine spectral}.
\newblock {\em Ann. Inst. Henri Poincar\'e}, {\bf 17}, (1981).
\newblock 31-49.

\bibitem{Wiener/Masani:57}
N.~Wiener and P.~Masani.
\newblock {The prediction theory of multivariable stochastic processes, I. The
  regularity condition}.
\newblock {\em Acta Math.}, {\bf 98}, (1957).
\newblock 111-150.

\bibitem{Wiener/Masani:58}
N.~Wiener and P.~Masani.
\newblock {The prediction theory of multivariable stochastic processes, II. The
  linear predictor}.
\newblock {\em Acta Math.}, {\bf 99}, (1958).
\newblock 93-137.

\end{thebibliography}

\begin{thebibliography}{10}

\bibitem{Blecher:95/2}
D.~P. Blecher.
\newblock {A generalization of Hilbert modules}.
\newblock preprint, University of Houston, Houston, Texas, USA, submitted to J.
  Funct. Anal., 1995.

\bibitem{Blecher:95/1}
D.~P. Blecher.
\newblock {A new approach to Hilbert C*-modules}.
\newblock preprint, University of Houston, Houston, Texas, USA, 1995.

\bibitem{Blecher:95/3}
D.~P. Blecher.
\newblock {On selfdual Hilbert modules}.
\newblock preprint, University of Houston, Houston, Texas, USA, submitted to J.
  Funct. Anal., 1995.

\bibitem{Blecher:96}
D.~P. Blecher.
\newblock {Matrix normed categories and Morita equivalence}.
\newblock preprint, Univ. Houston, Texas, U.S.A., 1996.

\bibitem{Blecher/Muhly/Na:96/1}
D.~P. Blecher, P.~S. Muhly, and Qiyuan Na.
\newblock {Morita equivalence of operator algebras and their C*-envelopes}.
\newblock preprint, 1996.

\bibitem{Blecher/Muhly/Paulsen:94}
D.~P. Blecher, P.~S. Muhly, and V.~I. Paulsen.
\newblock {Categories of operator modules (Morita equivalence and projective
  modules)}.
\newblock preprint, University of Iowa, Iowa City, USA, 1994.

\bibitem{Blecher/Muhly/Paulsen:96}
D.~P. Blecher, P.~S. Muhly, and V.~I. Paulsen.
\newblock {Categories of operator modules - Morita equivalence and projective
  modules}.
\newblock preprint, to appear in Memoirs Amer. Math. Soc., 1996.

\bibitem{Carlson/Clark:93}
J.~F. Carlson and D.~N. Clark.
\newblock {Ext for Hilbert modules}.
\newblock {\em Contemp. Math.}, {\bf 137}, (1992).
\newblock 113-118.

\bibitem{Carlson/Clark:95}
J.~F. Carlson and D.~N. Clark.
\newblock {Cohomology and extensions of Hilbert modules}.
\newblock {\em J. Functional Analysis}, {\bf 128}, (1995).
\newblock 278-306.

\bibitem{Carlson/Clark/Foias/Williams:94}
J.~F. Carlson, D.~N. Clark, C.~Foias, and J.~P. Williams.
\newblock {Projective Hilbert A(D)-modules}.
\newblock {\em New York J. Math.}, {\bf 1}, (1994).
\newblock 26-38, (Internet: http//nyjm.albany.edu or gopher://nyjm.albany.edu
  or ftp://nyjm.albany.edu).

\bibitem{Chen/Douglas:92}
Xiaoman Chen and R.~G. Douglas.
\newblock {Localization of Hilbert modules}.
\newblock {\em Mich. Math. J.}, {\bf 39}, (1992).
\newblock 443-454.

\bibitem{Douglas:86}
R.~G. Douglas.
\newblock {Hilbert modules over function algebras}.
\newblock In {\em {Advances in invariant subspaces and other results of
  operator theory, 9th Int. Conf. Oper. Theory, Timisoara and Herculane,
  Romania, 1984}}, volume~{\bf 17} of {\em {\it Oper. Theory, Adv. Appl.}},
  (1986).
\newblock 125-139.

\bibitem{Douglas:88}
R.~G. Douglas.
\newblock {On Silov resolution of Hilbert modules}.
\newblock In {\em {Special classes of linear operators and other topics, 11.
  Int. Conf. Oper. Theory, Bucharest, Romania, 1986}}, volume~{\bf 28} of {\em
  {\it Oper. Theory, Adv. Appl.}}, (1988).
\newblock 51-60.

\bibitem{Douglas:90}
R.~G. Douglas.
\newblock {Invariants for Hilbert modules}.
\newblock In {\em {Operator theory, operator algebras and applications, Proc.
  Summer Res. Inst., Durham/NH, U.S.A., 1988}}, volume {\bf 51}, Pt.~1 of {\em
  {\it Proc. Symp. Pure Math.}}, (1990).
\newblock 179-196.

\bibitem{Douglas:95}
R.~G. Douglas.
\newblock {Models and resolutions for Hilbert modules}.
\newblock {\em Contemp. Math.}, {\bf 185}, (1995).
\newblock 109-131.

\bibitem{Douglas/Paulsen:89}
R.~G. Douglas and V.~I. Paulsen.
\newblock {\em {Hilbert modules over function algebras}}.
\newblock Pitman Res. Notes Math. v.217. Longman, New York, 1989.

\bibitem{Douglas/Paulsen/Sah/Yan:95}
R.~G. Douglas, V.~I. Paulsen, C.~H. Sah, and K.~Yan.
\newblock {Algebraic reduction and rigidity for Hilbert modules}.
\newblock {\em Amer. J. Math.}, {\bf 117}, (1995).
\newblock 75-92.

\bibitem{Misra/Pati:93}
Gadadhar Misra and Vishwambhar Pati.
\newblock {Contractive and completely contractive modules, matricial tangent
  vectors and distance decreasing metrics}.
\newblock {\em J. Oper. Theory}, {\bf 30}, no.~2, (1993).
\newblock 353-380.

\bibitem{Muhly/Solel:93}
P.~S. Muhly and B.~Solel.
\newblock {Hilbert modules and partial orders}.
\newblock {\em C. R. Acad. Sci. Paris, Ser.~I}, {\bf 316}, (1993).
\newblock 1187-1191.

\bibitem{Muhly/Solel:95}
P.~S. Muhly and B.~Solel.
\newblock {Hilbert modules over operator algebras}.
\newblock {\em Memoirs Amer. Math. Soc.}, {\bf 559}, (1995).

\bibitem{Na:95}
Qiyuan Na.
\newblock {\em {Some contributions to the theory of operator modules}}.
\newblock PhD thesis, University of Iowa, Iowa City, U.S.A., 1995.

\bibitem{Sun/Li:88}
Shunhua Sun and Bingren Li.
\newblock {Synthetic techniques in operator theory}.
\newblock {\em Adv. Math., Beijing}, {\bf 17}, (1988).
\newblock 225-237.

\end{thebibliography}

\documentstyle{article}
\begin{document}

\begin{center}
{\large \bf Hilbert H*-modules}\\
Michael Frank \copyright \\
frank@mathematik.uni-leipzig.d400.de or \\
frank@rz.uni-leipzig.de
\end{center}

\noindent
\rule{6in}{0.1mm}

\noindent
last update: 1.4.1996

\nocite{Biedenharn/Horwitz:84,Biedenharn/Horwitz:85}
\nocite{Delanghe/Brackx:78,Delanghe:79,Giellis:72,Goldstine/Horwitz:66}
\nocite{Molnar:92,Molnar:90,Molnar:92/2,Molnar:92/3,Molnar:93/1,Molnar:93/2}
\nocite{Molnar:91,Razon/Horwitz:92,Razon/Horwitz:91/1,Razon/Horwitz:91/2}
\nocite{Cabrera/Martinez/Rodriguez:95,Cabrera/Martinez/Rodriguez:89/2}
\nocite{Saworotnow:68,Saworotnow:70,Saworotnow:85,Sharma/Coulson:87}
\nocite{Soffer/Horwitz:83,Smith:74,Cnops:92}
\nocite{Saworotnow:80,Saworotnow:82,Smith:73,Bultheel:82,Kulkarni:94}
\nocite{Saworotnow:91,Cabrera:91,Saworotnow:81,Wiener/Masani:57}
\nocite{Wiener/Masani:58,Wiener/Masani:60,Saworotnow:83,Kokschal:95,Fuge:95}
\nocite{Masani:59,Masani:62,Masani:66,Truong-Van:81,Salehi:65,Salehi:66}
\nocite{Salehi:67,Rosenberg:64,Ellis/Gohberg/Lay:95,Ben-Artzi/Gohberg:94}
\nocite{Sen:82,Bultheel:80,Cabrera/Martinez/Rodriguez:92}
\nocite{Cabrera/ElMarrakchi/Martinez/Rodriguez:93,Razon/Horwitz/Biedenharn:89}

\newpage

\begin{center}
{\large \bf Hilbert modules over operator algebras}  \\
Michael Frank \copyright \\
frank@mathematik.uni-leipzig.d400.de or \\
frank@rz.uni-leipzig.de
\end{center}

\noindent
\rule{6in}{0.1mm}

\noindent
last update: 1.4.96

\vspace{1cm}

\medskip \noindent
The collection of sources below is far from being complete, but
nevertheless, instructive as guide into the research field.

{\small

\nocite{Blecher/Muhly/Paulsen:96,Blecher:96,Na:95,Muhly/Solel:95}
\nocite{Douglas/Paulsen:89,Blecher:95/1,Blecher:95/2,Blecher:95/3}
\nocite{Blecher/Muhly/Na:96/1,Blecher/Muhly/Paulsen:94,Muhly/Solel:93}
\nocite{Carlson/Clark:95,Misra/Pati:93,Douglas/Paulsen/Sah/Yan:95}
\nocite{Douglas:95,Carlson/Clark/Foias/Williams:94,Chen/Douglas:92}
\nocite{Carlson/Clark:93,Douglas:90,Sun/Li:88,Douglas:88,Douglas:86}

}

\end{document}